# MonoDETRNext: Next-generation Accurate and Efficient Monocular 3D Object Detection Method

### v1

@@ 4.1.1 Dataset
MonoDETRNext is assessed on the extensively utilized KITTI[51] benchmark, which encompasses 7,481 training and 7,518 test images. Following the protocol outlined in previous studies [52, 53], we segregate 3,769 validation images from the training set. Evaluation is conducted across three difficulty levels: easy, moderate, and hard. Performance metrics are computed based on average precision ($AP$) of bounding boxes in both 3D space ($AP_{3D}$) and the bird’s-eye view ($AP_{BEV}$). These metrics are assessed at 40 recall positions to ensure comprehensive evaluation.
Note: The reason for not using commonly used datasets for other types of 3D detection methods is that the vast majority of monocular 3D detection methods have only been validated on the KITTI, and we cannot find their results on other datasets.
4.1.2 Implementation details
We employed ResNet50[54] as our backbone. While we also explored more advanced networks such as ConvNext[55], regrettably, these sophisticated architectures did not yield performance improvements. We utilized eight attention heads for all attention modules, including RGFI. MonoDETRNext-F was trained for 230 epochs on a single RTX 4090 GPU, with a batch size of 8 and a learning rate of $2\times 10^{-4}$. MonoDETRNext-A was trained for 300 epochs, with the learning rate decayed between epochs 225 and 275. We employed the AdamW[56] optimizer with weight decay of $10^{-4}$. For MonoDETRNext-F, the learning rate was decreased by a factor of 0.1 at epochs 145 and 205 and for MonoDETRNext-A, the process is between 165 and 205. The number of object queries was consistent with MonoDETR, set at 50. Inference speed was evaluated on a single RTX 3090 GPU.
4.2 Comparison with SOTA
@@
SSD-MonoDETR[68] - 24.52 17.88 15.73 33.59 24.35 21.98 29.53 21.96 18.20
MonoDETRNext-F 29.2 27.21 21.69 20.16 34.56 28.12 26.31 30.28 24.77 20.38
MonoDETRNext-A 20.8 29.94 24.14 23.79 37.32 30.68 30.29 33.44 25.75 22.40
@@
The table 1 presents a comparison of various monocular 3D object detection models on the KITTI dataset, focusing on the car category’s performance. The evaluation metrics $AP_{3D}$ and $AP_{BEV}$ , which are both at 40 recall positions, across different difficulty levels (Easy, Moderate, and Hard) for both the test and validation sets. Additionally, the FPS metric for batch size 1 is provided to gauge the inference speed of each model.
Among the methods listed, our methods MonoDETRNext-F and MonoDETRNext-A stands out. These variants exhibit notable advantages over other models in terms of both accuracy and efficiency.
@@
In terms of accuracy, MonoDETRNext-F achieves impressive $AP_{3D}$ and $AP_{BEV}$ scores across all difficulty levels, surpassing all other models listed. Specifically, it outperforms MonoDETR and SSD-MonoDETR in $AP_{3D}$, showcasing its superior ability to accurately detect 3D objects. Additionally, it achieves the highest $AP_{BEV}$ scores, indicating its effectiveness in detecting objects from a bird’s eye view perspective.
MonoDETRNext-A also demonstrates competitive performance, particularly in the validation set where it achieves the highest $AP_{3D}$ and $AP_{BEV}$ scores among all models.
Furthermore, both MonoDETRNext-F and MonoDETRNext-A exhibit respectable inference speeds, with MonoDETRNext-F achieving the highest $FPS_{bz=1}$ among all models.
@@ 4.3 Ablation Studies
Our study evaluates different strategies for object query generation, as depicted in Table 2(a). "L-center assign learnable embeddings," "enc-box," and "enc-center" refer to the anchors derived from backbone features, generated separately by two heads and contributing to individual loss calculation. Additionally, "enc-output" signifies six-dimensional anchors produced by a single head, also participating in loss calculation. Among these strategies, the proposed effective depth-aware query generation demonstrates notable performance, as evidenced by superior results across Easy, Moderate, and Hard stages compared to alternative methods.
Table 2(b) illustrates the performance of MonoDETRNext-F with different vision encoders, including the "hybrid" encoder featured in Figure 2. Notably, the "hybrid" encoder demonstrates superior performance across Easy, Moderate, and Hard categories compared to other configurations. Interestingly, even with fewer layers, the "hybrid-L" encoder maintains competitive performance, suggesting that a lower layer count may be more effective. This observation prompts speculation: the efficacy of fewer layers could stem from the relatively limited KITTI dataset or indicate that a single layer may constitute a more rational design choice. Further analysis is warranted to discern whether this trend holds across diverse datasets and to explore potential implications for architectural design in 3d object detection tasks.
In Table 2(d), the impact of different depth positional encodings on MonoDETRNext-F is examined. "3D sin/cos" and "2D sin/cos" utilize sinusoidal functions for encoding, with the former incorporating all six output dimensions. Notably, "3D sin/cos" outperforms other settings across Easy, Moderate, and Hard categories. This suggests that the precise encoding of depth information through sinusoidal functions significantly benefits object detection. Conversely, "Meter-wise" and "k-bin" positional encodings yield comparatively lower performance, indicating less effective depth representation in these configurations. Further investigation into the mechanisms behind these encoding methods is warranted to optimize depth-aware object detection algorithms.
In Table 2(c), $N_{1}$, $N_{2}$, and $N_{3}$ signify the number of SDC modules, as depicted in Figure 2 .
(a) Performance of MonoDETRNext-F at different 3D object query generation strategy
@@
L-center+enc-box 30.28 24.77 20.38
enc-center+enc-box 23.89 19.49 17.97
enc-output 23.80 18.96 16.47
w/o 22.97 17.01 15.45
(b) Performance of MonoDETRNext-F under different vision encoder
@@
(c) Effect of Different Scales of Accurate Depth Predictors on MonoDETRNext-A Performance
@@
4, 4, 6 31.78 23.56 20.78
(d) The impact of different depth positional encodings of MonoDETRNext-F.
@@
Table 2: Ablation Studies on Our Proposed MonoDETRNext on the KITTI validation benchmark. The specific meanings of the table can be referred to in Section 4.3.
@@ 5 Conclusion and Limitations
Conclusion: This paper introduces a novel monocular vision-based approach for 3D object detection. Leveraging advancements from the 2D detection domain, we propose the efficient and precise MonoDETRNext. Building upon the groundwork laid by MonoDETR, we introduce two variants: MonoDETRNext-F prioritizing speed, and MonoDETRNext-A emphasizing accuracy. Our methodology encompasses the development of an efficient hybrid vision encoder, enhancements to depth prediction mechanisms, and refinements in object query generation. Through comprehensive performance evaluation, we establish the superiority of our model over existing methods. By optimizing both accuracy and computational efficiency, MonoDETRNext sets a new benchmark in monocular 3D object detection, facilitating future research and applications across diverse real-world scenarios.

### v2

@@ 4.1.1 Dataset
MonoCD [69] 31.7 25.53 16.59 14.53 33.41 22.81 19.57 26.45 19.37 16.38
DA3D [25] - 27.76 20.47 17.90 36.83 26.92 23.41 - - -
MonoDETRNext-E 29.2 27.35 21.24 19.16 34.73 27.32 24.87 30.28 24.77 20.38
MonoDETRNext-A 20.8 28.52 22.81 21.08 35.86 28.96 27.58 32.95 25.01 21.92
KITTI[16] encompasses 7,481 training and 7,518 test images. Following the protocol outlined in previous studies [11, 10], we segregate 3,769 validation images from the training set. Evaluation is conducted across three difficulty levels: easy, moderate, and hard. Performance metrics are computed based on average precision ($AP$) of bounding boxes in both 3D space ($AP_{3D}$) and the bird’s-eye view ($AP_{BEV}$). These metrics are assessed at 40 recall positions to ensure comprehensive evaluation.
[…]
4.1.2 End-to-End 3D object tracking
To further demonstrate versatility, we also extend MonoDETRNext to 3D Multi-Object Tracking (MOT), the specific tracking approach is consistent with MOTR [75], interested readers can read it by themselves, here we use the same training method as [52]. Through this experiment, we demonstrate that the query-based MOT model has higher accuracy in 3d tracking when using the 3d detection model as the base model. The primary metrics employed for evaluation included HOTA, AssA and MOTA [56, 46]. These metrics collectively provided a comprehensive and robust evaluation of our model’s tracking performance.
4.1.3 Implementation details
We employed ResNet50[18] as our backbone. We utilized eight attention heads for all attention modules, including RGFI. MonoDETRNext-E was trained for 230 epochs on a single RTX 4090 GPU, with a batch size of 8 and a learning rate of $2\times 10^{-4}$. MonoDETRNext-A was trained for 300 epochs. We employed the AdamW [44] optimizer with weight decay of $10^{-4}$. For MonoDETRNext-E, the learning rate was decreased by a factor of 0.1 at epochs 145 and 205 and for MonoDETRNext-A, the process is between 165 and 205. For the Rope3D, with 150 epochs of both the MonoDETRNext-E and A trained, the learning rate decreases starting at 80 epochs and ending at 120 epochs. Tracking, on the other hand, was trained for another 200 epochs on the MOT part of the KITTI dataset using the E version on top of the detection The number of object queries was consistent with MonoDETR, set at 50. Inference speed was evaluated on a single RTX 3090 GPU.
4.2 Main Results
@@
KITTI: The table 1 presents a comparison of various monocular 3D object detection models on the KITTI dataset, focusing on the car category’s performance. The evaluation metrics $AP_{3D}$ and $AP_{BEV}$ , which are both at 40 recall positions, across different difficulty levels (Easy, Moderate, and Hard) for both the test and validation sets. Additionally, the FPS metric for batch size 1 is provided to gauge the inference speed of each model.
@@
Among the methods listed, our methods MonoDETRNext-E and MonoDETRNext-A stands out. These variants exhibit notable advantages over other models in terms of both accuracy and efficiency. In terms of accuracy, MonoDETRNext-A achieves impressive $AP_{3D}$ and $AP_{BEV}$ scores across all difficulty levels, surpassing all other models listed. Specifically, it outperforms MonoDETR in $AP_{3D}$, showcasing its superior ability to accurately detect 3D objects. Additionally, it achieves the highest $AP_{BEV}$ scores, indicating its effectiveness in detecting objects from a bird’s eye view perspective. At the same time, we found that our method significantly outperforms other methods in both moderate and hard scenarios, which is probably because the introduction of anchors significantly improves the impact of occlusion on detection performance. Furthermore, MonoDETRNext-E exhibits respectable inference speeds, with MonoDETRNext-E achieving the second $FPS_{bz=1}$ among all models.
Rope3D: For the Rope3D performance comparison in Table 2, MonoDETRNext achieves state of-the-art results, with the MonoDETRNext-A variant surpassing all other methods, including BEVDepth and MonoFlex, in both 3D object detection ($AP_{3D}$ of 82.34) and $Rope_{score}$ (85.16). This highlights the model’s superiority in 3D scene understanding from monocular images.
Table 2: Performance comparison between MonoDETRNext and existing methods on the Rope3D [74] set.
Method $AP_{3D}$ $Rope_{score}$
MonoFlex[80] 59.78 66.66
GUPNet[45] 66.52 70.14
BEVDepth[32] 69.63 74.70
BEVHeight[72] 74.60 78.72
MonoDLE [48] 77.50 80.84
MonoDETR[79] 78.68 81.93
MonoDETRNext-E 81.89 83.69
MonoDETRNext-A 82.34 85.16
Tracking: As shown in Table 3, the tracking results indicate that MonoDETRNext-E-MOT sets a new benchmark on the KITTI, outperforming other end-to-end and non-end-to-end methods like CenterTr and PolarMOT. It achieves high HOTA, MOTA, and AssA, as well as relatively low IDSW, indicating its robustness and accuracy in target tracking. This suggests MonoDETRNext’s strong potential in both object detection and tracking tasks, offering a unified and efficient solution across multiple benchmarks.
Table 3: The tracking performance on KITTI after converting MonoDETRNext into an end-to-end tracking model.
[…]
Non-End-to-end:
CenterTr [83] 73.02 88.83 71.12 75.62 254
PolarMOT [27] 75.16 85.08 76.95 73.94 462
TripletTrack [50] 73.58 84.32 74.66 73.18 322
PermaTr [59] 78.03 90.13 78.41 78.29 258
End-to-end:
DecoderTracker[52] 72.08 83.19 73.84 71.84 275
MonoDETRNext-E-MOT 78.93 90.27 81.11 73.05 143
@@ 4.3 Ablation Studies
Our study evaluates different strategies for object query generation, as depicted in Table 4. ”L-center assign learnable embeddings,” ”enc-box,” and ”enc-center” refer to the anchors derived from backbone features, generated separately by two heads and contributing to individual loss calculation. From the table, it is evident that MonoDETRNext significantly improves detection performance in moderate and challenging conditions. This improvement suggests that the strategy of using queries to generate anchors can mitigate some of the adverse effects caused by object occlusions.
Table 4: Performance of MonoDETRNext-E at different 3D object query generation strategy.
@@
L-center 25.80 18.96 17.47
L-center+enc-box 27.89 21.49 19.96
enc-center 28.89 20.78 18.52
enc-center+enc-box 30.28 24.77 20.38
Table 5 illustrates the performance of MonoDETRNext-E with different vision encoders, including the ”hybrid” encoder featured in Figure 2. Notably, the ”hybrid” encoder demonstrates superior performance across Easy, Moderate, and Hard categories compared to other configurations. Interestingly, even with fewer layers, the ”hybrid-L” encoder maintains competitive performance, suggesting that a lower layer count may be more effective.
Table 5: Performance of MonoDETRNext-E under different vision encoder.
@@
MonoDETR’s 28.35 22.53 18.74
hybrid-L 29.67 23.81 19.03
our hybrid 30.28 24.77 20.38
In Table 6, $N_{1}$, $N_{2}$, and $N_{3}$ signify the number of SDC modules, as depicted in Figure 2. The configuration $(2,2,4)$ yields the best results across all levels, indicating an optimal balance. Both smaller $(1,1,2)$ and larger scales $(3,3,6)$ and $(4,4,6)$ show diminished performance, suggesting issues with underfitting and overfitting respectively.
Table 6: Effect of Different Scales of Accurate Depth Predictors on MonoDETRNext-A Performance.
@@
4, 4, 6 31.78 23.92 20.78
The Table.7 presents an ablation study on the depth estimation network module in MonoDETRNext-A. Removing the point-wise operations improves performance across all difficulty levels, particularly evident in the ”Easy” category, highlighting their significant contribution to depth estimation accuracy.
Table 7: The impact of removing point-wise in the depth estimation network module on the performance of MonoDETRNext-A. ✓indicates deletion, and ✗indicates no deletion.
SDC RGFI Easy Moderate Hard
29.62 22.43 18.86
31.86 22.86 22.04
31.54 23.42 21.81
32.95 25.01 21.92
In Table 8, the impact of different depth positional encodings on MonoDETRNext-E is examined. ”3D sin/cos” and ”2D sin/cos” utilize sinusoidal functions for encoding, with the former incorporating all six output dimensions. Notably, ”3D sin/cos” outperforms other settings across Easy, Moderate, and Hard categories. This suggests that the precise encoding of depth information through sinusoidal functions significantly benefits object detection. Conversely, ”Meter-wise” and ”k-bin” positional encodings yield comparatively lower performance, indicating less effective depth representation in these configurations.
Table 8: The impact of different depth positional embeddings of MonoDETRNext-E.
@@ 5 Conclusion and Limitations
Conclusion: This paper introduces a novel monocular vision-based approach for 3D object detection. Leveraging advancements from the 2D detection domain, we propose the efficient and precise MonoDETRNext. Building upon the groundwork laid by MonoDETR, we introduce two variants: MonoDETRNext-E prioritizing speed, and MonoDETRNext-A emphasizing accuracy. Our methodology encompasses the development of an efficient hybrid vision encoder, enhancements to depth prediction mechanisms, and refinements in object query generation. Through comprehensive performance evaluation, we establish the superiority of our model over existing methods. By optimizing both accuracy and computational efficiency, MonoDETRNext sets a new benchmark in monocular 3D object detection, facilitating future research and applications across diverse real-world scenarios.
@@
Supplementary Material
Figure 5: The visualization results of the final attention output from the decoder of MonoDETR and MonoDETRNext. Hotter colors indicate higher attention weights. It is evident that the heat map of MonoDETRNext is more concentrated than that of MonoDETR.
6 Some discussions on the success of MonoDETRNext
The success of MonoDETRNext can be attributed to a comprehensive understanding of the limitations inherent in existing models, particularly MonoDETR, and the innovative strategies employed to address these challenges. By critically analyzing the shortcomings of MonoDETR, the authors identified key areas for improvement, leading to significant advancements in monocular 3D object detection.
One of the most notable enhancements in MonoDETRNext is the optimization of query initialization. In the original MonoDETR framework, query initialization posed constraints that adversely affected detection efficiency and accuracy. MonoDETRNext tackles this issue by implementing a refined initialization strategy that allows the model to more effectively capture crucial scene features, thus enhancing overall detection performance. This improvement demonstrates the importance of fine-tuning initial conditions to bolster model efficacy in dynamic environments.
The introduction of a more powerful depth estimator is another cornerstone of MonoDETRNext’s advancements. This enhanced estimator significantly increases accuracy in depth perception, allowing the model to construct a more precise understanding of three-dimensional structures within a scene. By leveraging advanced deep learning techniques, the new depth estimator minimizes errors in complex environments, contributing to higher overall detection quality.
In Figure 5, we present the visualization heatmaps of the attention outputs from our two models and MonoDETR. A closer examination reveals that our models exhibit a more concentrated and accurate attention distribution in key areas. This indicates that our models are more effective in identifying and focusing on critical information within the images, thereby enhancing detection accuracy.
In conclusion, the achievements of MonoDETRNext underscore the significance of addressing fundamental issues within existing frameworks. Through innovative enhancements in query initialization, encoder architecture, and depth estimation, the model sets a new standard for monocular 3D object detection. The accompanying heatmap analyses provide compelling evidence of these advancements, affirming the model’s potential to influence future research and applications in the field.
7 How to achieve better detection accuracy on KITTI
It is well known that one of the best ways to enhance the performance of deep learning models is by increasing the dataset size. In pursuit of maximizing the performance of MonoDETRNext, we attempted to expand the dataset.
The original Repo3D dataset was a reasonable choice; however, its primary focus on traffic signals does not align well with the scenarios found in the KITTI dataset. Therefore, we turned our attention to another dataset commonly used in autonomous driving, nuScenes. The nuScenes dataset contains a wealth of driving scenarios, encompassing urban streets, rural roads, and various weather conditions, making it highly suitable for training and evaluating 3D object detection models related to autonomous driving.
The nuScenes dataset comprises data from six cameras, and we selected the forward camera as our supplementary data source. Additionally, since the nuScenes dataset does not allow for the calculation of the $\alpha$ angle when converting to KITTI format, we only utilized it to supervise the depth module. One more thing ,similar to LPCG[53], we generate pseudo labels using KITTI raw data (42K unlabeled data) to add additional training data, further improving the performance of our method. Consequently, we focused our efforts on MonoDETRNext-A. The results are presented in Table 9.
Table 9: Monocular performance of the car category on KITTI test and val sets. The top three results are highlighted in the order of red, green, and blue.
Method Test, $AP_{3D}$ Test, $AP_{BEV}$
Easy Mod. Hard Easy Mod. Hard
MonoPair [13] 13.04 9.99 8.65 19.28 14.83 12.89
DDMP-3D [61] 19.71 12.78 9.80 28.08 17.89 13.44
MonoRUn [8] 19.65 12.30 10.58 27.94 17.34 15.24
CaDDN [55] 19.17 13.41 11.46 27.94 18.91 17.19
MonoDLE [48] 17.23 12.26 10.29 24.79 18.89 16.00
MonoRCNN [58] 18.36 12.65 10.03 25.48 18.11 14.10
MonoGeo [57] 18.85 13.81 11.52 25.86 18.99 16.19
GUPNet [45] 20.11 14.20 11.77 - - -
MonoDTR [22] 21.99 15.39 12.73 28.59 20.38 17.14
AutoShape [42] 22.47 14.17 11.36 30.66 20.08 15.59
MonoDETR[79] 25.00 16.47 13.58 33.60 22.11 18.60
SSD-MonoDETR[19] 24.52 17.88 15.73 33.59 24.35 21.98
MonoTAKDv2 [36] 29.86 21.26 18.27 43.83 32.31 28.48
DA3D* [25] 30.83 22.08 19.20 39.50 28.71 25.20
MonoDETRNext-E 27.35 21.24 19.16 34.73 27.32 24.87
MonoDETRNext-A 28.52 22.81 21.08 35.86 28.96 27.58
MonoDETRNext-A(Extra Data) 29.94 24.14 23.79 37.32 30.68 30.29